OpenStudy (anonymous):

What is the sum of the first five terms of a geometric series with a sub 1 = 6 and r = 1/3? Express your answer as an improper fraction in lowest terms without using spaces.

OpenStudy (campbell_st):

ok so the sum of a geometric series is \[s_{5} = \frac{6 \times (1 - (\frac{1}{3})^5}{1 - \frac{1}{3}}\] just evaluate it

OpenStudy (anonymous):

that looks so complicated @campbell_st

OpenStudy (anonymous):

@campbell_st 322 over 81?

OpenStudy (campbell_st):

I got 242/27